OpenStudy (anonymous):

Plss i need solution or maybe an answer i just want to make sure if my solutions are right here the question its Calculus >>>f(x)''''' of f(x)=15x^4-(2x-1/2)^6

OpenStudy (misty1212):

lol hi!

OpenStudy (misty1212):

is that the fifth derivative?

OpenStudy (anonymous):

\[ f(x)=15x ^{4}-(\frac{ 2x-1 }{ 2 })^{6}\]

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

i need some solutions or answers already :)

OpenStudy (kainui):

Do you know how to take the first derivative?

OpenStudy (anonymous):

Yup i already got the solutions :) i just want to compare if my answers correct

OpenStudy (anonymous):

the degree of the first term is 4 so it will disappear in the fifth derivative.

OpenStudy (anonymous):

Yeah that my answer is it right?

OpenStudy (anonymous):

-360(2x-1)

OpenStudy (anonymous):

the second terms will leave this after 5 derivatives

OpenStudy (anonymous):

then its -720x+360 right ?

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

the sixth derivative leaves -720 and seventh derivative leaves 0

OpenStudy (anonymous):

That professor !!!! -_- she said im wrong :(

OpenStudy (anonymous):

Our teacher said that my answers wrong?

OpenStudy (anonymous):

\[f'(x) =15\times 4\times x^3 - 6(x-\frac{1}{2})^5\] \[f''(x) =15\times4\times 3 x^2 - 6\times5(x-\frac{1}{2})^4\]

OpenStudy (anonymous):

Umm.. Its ok i know already how to get it i just want know if im wrong .. ^^ Thanks

OpenStudy (anonymous):

\[f'''(x) =15\times4\times3 \times2x - 6\times5\times4(x-\frac{1}{2})^3\] \[f''''(x) =15\times4\times3 \times2\times1 - 6\times5\times4\times3(x-\frac{1}{2})^2\]\[f'''''(x) =15\times4\times3 \times2\times1\times0 - 6\times5\times4\times3\times2(x-\frac{1}{2})^1\] =0 - 720x+360 Your answer is absolutely correct. (according to leibnitz)

OpenStudy (anonymous):

:)